If a,b,c,d are proportional, prove that a+d=b+c+(a-b)(a-c)                                                                                     a

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
If a,b,c,d are proportional, To prove
a+d=b+c+((a-b)(a-c)/a)
a(a-b-c+d) = a^2-ab+ad=a^2 - ab - ac +bc = (a-b)(a-c);
a-b-c+d = +(a-b)(a-c)/a
Soa+d=b+c+((a-b)(a-c)/a)